XyMTeX Packages for Chemists v4.06

Shinsaku Fujita Shonan Institute of Chemoinformatics and Mathematical Chemistry

Kaneko 479-7 Ooimachi, Ashigara-Kami-Gun, Kanagawa, 258-0019 Japan

2009/12/01

Contents

1 docstrip options

This dtx file provides the xymtex package by docstrip processing.

argument package name
xymtex  xymtex.sty
xymtexps xymtexps.sty
driver  a documentation driver file

2 Version information

The old process for typing out version information remains active.

 1 ⟨xymtex⟩\def\j@urnalname{xymtex} \def\journalID{xymtex}
 2 ⟨xymtexps⟩\def\j@urnalname{xymtexps} \def\journalID{xymtexps}
 3 \def\versi@ndate{December 01, 2009}
 4 \def\versi@nno{ver4.06}
 5 \def\copyrighth@lder{SF} % Shinsaku Fujita
 6 \typeout{Package `\j@urnalname' (\versi@nno) <\versi@ndate>\space
 7 [\copyrighth@lder]}
 8 ⟨xymtex⟩ \typeout{xymtex.sty ver4.06 (for LaTeX2e) 2009/12/01[SF]}
 9 ⟨xymtexps⟩ \typeout{xymtexps.sty ver4.06 (for LaTeX2e) 2009/12/01[SF]}
10 \usepackage{epic}
11 \usepackage{chemstr}
12 \usepackage{carom}
13 \usepackage{hetarom}
14 \usepackage{hetaromh}
15 \usepackage{lowcycle}
16 \usepackage{ccycle}
17 \usepackage{hcycle}
18 \usepackage{aliphat}
19 \usepackage{locant}
20 \usepackage{polymers}
21 \usepackage{methylen}
22 \usepackage{fusering}

23 \usepackage{sizeredc}
24 \usepackage{steroid}
25 \usepackage{lewisstruc}
26 ⟨xymtexps⟩\usepackage{xymtx-ps}